ow high should a 3kg mass be lifted from the ground if it is thrown upward at a speed of 25m/s
(Take g=10m/s?
1
Expert's answer
2019-03-11T12:19:14-0400

If there is no air friction, apply law of conservation of energy: the kinetic energy will provide the potential energy (i.e. lifting above the ground):


"KE=PE,"

"\frac{1}{2}mv^2=mgh,"

"h=\frac{v^2}{2g}=31.25 \text{ m}."

The height does not depend on mass.